inventory information for Part 311 of Tamarisk Corp. discloses the following information for the month of June.

June   1

Balance

302 units @ $15

June 10

Sold

200 units @ $35

11

Purchased

796 units @ $18

15

Sold

499 units @ $37

20

Purchased

501 units @ $19

27

Sold

301 units @ $40

(a)

Assuming that the periodic inventory method is used, compute the cost of goods sold and ending inventory under (1) LIFO and (2) FIFO.

(1)
LIFO

(2)
FIFO

Cost of Goods Sold

$enter a dollar amount

$enter a dollar amount

Ending Inventory
Answer #1
Unit Unit Cost Total Cost
Beg Inv 302 15 4530
Purchase 1 796 18 14328
Purchase 2 501 19 9519
Total 1599 28377

Sales unit = 200+499+301 = 1000 Units

Calculate following

LIFO FIFO
Cost of goods sold 9519+499*18 = 18501 4530+698*18 = 17094
Ending inventory 28377-18501 = 9876 28377-17094 = 11283